OpenStudy (anonymous):

FAN/MEDAL!!! For the graphed function f(x) = −(5)x − 3 + 2, calculate the average rate of change from x = 4 to x = 6

OpenStudy (anonymous):

answer choices are: −68 −60 68 60

OpenStudy (anonymous):

take ist derivaite first

OpenStudy (anonymous):

I dont know how to do that? @Nurali

OpenStudy (anonymous):

@Brainybeauty

OpenStudy (anonymous):

\[f(x) = −(5)^{x − 3} + 2\] maybe?

OpenStudy (anonymous):

can u state the question in correct form plz

OpenStudy (anonymous):

Thats exactly how it is written on my flvs

OpenStudy (anonymous):

whatever it is, the answer is \[\frac{f(6)-f(4)}{6-4}\]

OpenStudy (anonymous):

@cambam1323 the way you wrote the function is not clear

OpenStudy (anonymous):

how do i solve that? do i have to do something special to the f(^6) and f(4)? i really need help!

OpenStudy (anonymous):

but whatever the function is, your job is to compute \[\frac{f(6)-f(4)}{2}\]

OpenStudy (anonymous):

since i cannot read \(f(x)\) i cannot figure out what \(f(6)\) is

OpenStudy (anonymous):

this is exactly what it looks like... For the graphed function f(x) = −(5)x − 3 + 2, calculate the average rate of change from x = 4 to x = 6. graph of f of x equals negative 1 times 5 to the x minus 3 power, plus 2 −68 −60 68 60

OpenStudy (anonymous):

it is the function that is illegible it is \(f(x) = −(5)^{x − 3} + 2\)? \[ f(x) = −(5)x − 3 + 2\] makes no sense

OpenStudy (anonymous):

\[f(x)=-(5)^{x-3}+2\]

OpenStudy (anonymous):

sorry about that

OpenStudy (anonymous):

yay

OpenStudy (anonymous):

\[f(6)=-5^{6-3}+2=-5^3+2=-125+2=-123\]

OpenStudy (anonymous):

your turn for \(f(4)\)

OpenStudy (anonymous):

I dont know how you got that? care to explain so i can then try?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[\large f(x)=-5^{x-3}+2\\ f(\spadesuit)=-5^{\spadesuit-3}+2\\ f(\heartsuit)=-5^{\heartsuit-3}+2\\ f(\clubsuit)=-5^{\clubsuit-3}+2\\ f(\diamondsuit)=-5^{\diamondsuit-3}+2\\ f(6)=-5^{6-3}+2\]

OpenStudy (anonymous):

in english , where you see an \(x\) replace it by a \(6\)

OpenStudy (anonymous):

and so \[f(4)\] where you see a \(x\) replace it by \(4\) it is how you evaluate a function

OpenStudy (anonymous):

okay so it would be \[f(4)=-5^{4-3}+2\]

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

although you still have to compute that number

OpenStudy (anonymous):

how just through order of opperations?

OpenStudy (anonymous):

i guess so

OpenStudy (anonymous):

so f(4)=7???

OpenStudy (anonymous):

i get minus three

OpenStudy (anonymous):

oh i see you forgot the minus sign in front of the five

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

\[f(4)=-5^{4-3}+2=-5^1+2=-5+2=-3\]

OpenStudy (anonymous):

i think you dropped the minus sign

OpenStudy (anonymous):

ohh yeah i did. :(

OpenStudy (anonymous):

ok now we are almost done

OpenStudy (anonymous):

yay!

OpenStudy (anonymous):

\[f(6)=-123\\ f(4)=-3\] and you need to compute \[\frac{f(6)-f(4)}{6-4}\]

OpenStudy (anonymous):

plug in the numbers, be careful with the arithmetic

OpenStudy (anonymous):

\[\frac{ -123-(-3) }{6-4 }\]

OpenStudy (anonymous):

yes ...

OpenStudy (anonymous):

and then i just solve it?

OpenStudy (anonymous):

let me know when you get \(-60\)

OpenStudy (anonymous):

if by "solve" you mean "compute the number" then yes

OpenStudy (anonymous):

i got -60 is that correct?!? :)

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

hooray!!! thank you so so so so so so so much!!!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

i an do some math, can play no instruments though...

OpenStudy (anonymous):

haha i spend most of my time playing and learning instruments and reading, i try to get ahead in school too. i also sing in a band. and i do lots of sports so im pretty busy :)